OpenStudy (anonymous):

solve the equation: ln square root (x+7)=4 please show me step by step:)... thanks

OpenStudy (anonymous):

\[\sqrt{x+7}=4\]?

OpenStudy (anonymous):

squaring both sides, we get x+7 = 16 x = 16-7 x = 9

OpenStudy (anonymous):

square both sides so you get +/-(x+7)=16 then solve for both cases: case one x+7=16 x=9 case two: -(x+7)=16 x+7=-16 x=-23 check for extraneous solutions and we find that if x=-23 we will get sqareroot of a negative so x cannot be -23 thus x=9

OpenStudy (anonymous):

\[ln \sqrt{x+7}=4\]

OpenStudy (anonymous):

natural log??

OpenStudy (anonymous):

dang... lol ok so then if you rewrite the natural log in exponenital form you get: sqrt(x+7)=e^4 then square both sides so x+7=e^8 x=(e^8)-7

OpenStudy (anonymous):

there is the negative case senario which would work as well where: -(x+7)=e^8 so x+7=-(e^8) thus x=(-(e^8))-7

OpenStudy (anonymous):

but that would give you an unreal answer which is not an option

OpenStudy (anonymous):

\[\frac{1}{2}\ln(x+7)=4\]\[\ln(x+7)=4*\frac{2}{1}\]\[\ln(x+7)=8\]\[x=e^8-7\]